# On the area of ordinary hyperbolic reduced polygons

### v1

@@ Abstract.
A convex body $R$ in the hyperbolic plane is reduced if any convex body $K\subset R$ has a smaller minimal width than $R$. We examine the area of a family of hyperbolic reduced $n$-gons, and prove that, within this family, regular $n$-gons have minimal area.
@@ 1. Introduction
On the hyperbolic plane among ordinary reduced $n$-gons of minimal width $w$, regular $n$-gons of the same minimal width have the smallest area.
@@ 3. Ordinary reduced polygons
As a consequence, we can derive the following area formula. Let $F_{w}\left(x\right)=f_{w}\left(g_{w}\left(x\right)\right)$, where
@@
$\cos\varphi_{i}=\frac{\tanh b_{i}}{\tanh c_{i}}=\frac{\tanh b_{i}}{\tanh\left(w-b_{i}\right)}=\frac{\tanh b_{i}\left(1-\tanh w\tanh b_{i}\right)}{\tanh w-\tanh b_{i}}.$
@@
$\tanh b_{i}=\frac{1+\cos\varphi_{i}-\sqrt{\left(1+\cos\varphi_{i}\right)^{2}-4\tanh^{2}w\cos\varphi_{i}}}{2\tanh w}=g_{w}\left(\varphi_{i}\right)$
@@
$\sinh\left(\operatorname{\mathrm{artanh}}\left(x\right)\right)=\frac{x}{\sqrt{1-x^{2}}},$
@@
$\sinh c_{i}=\sinh\left(\operatorname{\mathrm{artanh}}\left(\frac{\tanh w-\tanh b_{i}}{1-\tanh w\tanh b_{i}}\right)\right)=\sinh\left(\operatorname{\mathrm{artanh}}\left(\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}\right)\right)=$
$=\frac{\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}}{\sqrt{1-\left(\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}\right)^{2}}}=\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{\sqrt{\left(1-\tanh^{2}w\right)\left(1-g_{w}\left(\varphi_{i}\right)^{2}\right)}},$
@@
Now we can show that contrary with the Euclidean and spherical case, regular $n$-gons minimize the area among ordinary reduced $n$-gons.
@@
$\operatorname{\mathrm{area}}\left(R\right)\geq\operatorname{\mathrm{area}}\left(\widetilde{R}\right)$
@@
$r_{w}\left(x\right)=\sqrt{\left(1+\cos\varphi_{i}\right)^{2}-4\tanh^{2}w\cos\varphi_{i}}.$
We can calculate that the second derivative is
$F_{w}^{\prime\prime}\left(x\right)=\frac{\sqrt{1-\tanh^{2}w}}{\left(\tanh w\right)\left(\tanh^{2}w\left(g_{w}\left(x\right)+1\right)-2\tanh wg_{w}\left(x\right)\right)r_{w}^{2}\left(x\right)}h_{w}\left(x\right),$
@@
$h_{w}\left(x\right)=\sin x\left(\left(\tanh w-g_{w}\left(x\right)+1\right)\left(\sin x-r_{w}^{\prime}\left(x\right)\right)-g_{w}^{\prime}\left(x\right)\left(2\tanh^{2}w-r_{w}\left(x\right)-\cos x-1\right)\right)$
$-\cos x\left(\tanh w-g_{w}\left(x\right)+1\right).$
First we observe that
$\frac{\sqrt{1-\tanh^{2}w}}{\left(\tanh w\right)\left(\tanh^{2}w\left(g_{w}\left(x\right)+1\right)-2\tanh wg_{w}\left(x\right)\right)r_{w}^{2}\left(x\right)}>0,$
so the concavity of $F$ only depends on the sign of $h_{w}$. We also notice that for $0<x<\max_{i}\varphi_{i}<\frac{\pi}{2}$ both $\sin x$ and $\cos x$ are positive, $\tan w>g_{w}\left(x\right)$ by the argument in the proof of Theorem 3.4, and then we also have the following:
$\sin x-r_{w}^{\prime}\left(x\right)=\frac{\sin x}{r_{w}\left(x\right)}\left(r_{w}\left(x\right)+1+\cos\left(x\right)-2\tanh^{2}w\right)=\frac{2\tanh w\sin x}{r_{w}\left(x\right)}\left(g_{w}\left(x\right)-\tanh w\right)<0.$
On the other hand,
$g^{\prime}\left(x\right)=\frac{-\sin x+r_{w}^{\prime}\left(x\right)}{2\tanh w}=-\frac{\sin x}{r_{w}\left(x\right)}\left(g_{w}\left(x\right)-\tanh w\right),$
so
$h_{w}\left(x\right)=\frac{2\sin^{2}x\tanh w}{r_{w}\left(x\right)}\left(g_{w}\left(x\right)-\tanh w\right)\left(2\tanh w-2g_{w}\left(x\right)+1\right)-\cos x\left(\tanh w-g_{w}\left(x\right)+1\right),$
which is negative, and hence $-F_{w}$ is strictly convex in the interval $\left(0,\max_{i}\varphi_{i}\right]$. Finally, Jensen’s inequality implies the statement of the theorem. ∎
Since the limit of the areas of regular $n$-gons is the area of the circle of the same width as $n$ tends to infinity, this might sound like a contradiction if we consider Lassak’s conjecture about maximizing the area of reduced convex bodies. He conjectures that on the Euclidean plane, the circle and the quarter of the disk have maximal area. However, on one hand, hyperbolic width function does not tend to the Euclidean one as $w$ approaches 0, on the other hand, the quarter of the disk has a strictly greater area than the circle of the same thickness in $H^{2}$.

### v2

@@ Abstract.
A convex body $R$ in the hyperbolic plane is reduced if any convex body $K\subset R$ has a smaller minimal width than $R$. We examine the area of a family of hyperbolic reduced $n$-gons, and prove that, within this family, regular $n$-gons have maximal area.
@@ 1. Introduction
On the hyperbolic plane among ordinary reduced $n$-gons of minimal width $w$, regular $n$-gons of the same minimal width have the greatest area.
@@ 3. Ordinary reduced polygons
4. Area of ordinary reduced polygons
As a consequence of Lemma 3.3, we can derive the following area formula. Let $F_{w}\left(x\right)=f_{w}\left(g_{w}\left(x\right)\right)$, where
@@ 4. Area of ordinary reduced polygons
(1) $\cos\varphi_{i}=\frac{\tanh b_{i}}{\tanh c_{i}}=\frac{\tanh b_{i}}{\tanh\left(w-b_{i}\right)}=\frac{\tanh b_{i}\left(1-\tanh w\tanh b_{i}\right)}{\tanh w-\tanh b_{i}}.$
@@
(2) $\tanh b_{i}=\frac{1+\cos\varphi_{i}-\sqrt{\left(1+\cos\varphi_{i}\right)^{2}-4\tanh^{2}w\cos\varphi_{i}}}{2\tanh w}=g_{w}\left(\varphi_{i}\right)$
@@
$\sinh\left(\operatorname{\mathrm{artanh}}x\right)=\frac{x}{\sqrt{1-x^{2}}},$
@@
(3) $\sinh c_{i}=\sinh\left(\operatorname{\mathrm{artanh}}\left(\frac{\tanh w-\tanh b_{i}}{1-\tanh w\tanh b_{i}}\right)\right)=\sinh\left(\operatorname{\mathrm{artanh}}\left(\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}\right)\right)=$
(4) $=\frac{\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}}{\sqrt{1-\left(\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{1-\tanh w\cdot g_{w}\left(\varphi_{i}\right)}\right)^{2}}}=\frac{\tanh w-g_{w}\left(\varphi_{i}\right)}{\sqrt{\left(1-\tanh^{2}w\right)\left(1-g_{w}\left(\varphi_{i}\right)^{2}\right)}},$
@@
Now we are ready to prove our main result.
@@
$\operatorname{\mathrm{area}}\left(R\right)\leq\operatorname{\mathrm{area}}\left(\widetilde{R}\right)$
@@
$r_{w}\left(x\right)=\sqrt{\left(1+\cos x\right)^{2}-4\tanh^{2}w\cos x}.$
After differentiating, we get
$F_{w}^{\prime}\left(x\right)=\frac{g_{w}^{\prime}\left(x\right)\sqrt{1-\tanh^{2}w}\left(\tanh w-g_{w}\left(x\right)\right)+g_{w}\left(x\right)\sqrt{1-\tanh^{2}w}g_{w}^{\prime}\left(x\right)}{\sqrt{1-\left(\frac{g_{w}\left(x\right)\sqrt{1-\tanh^{2}w}}{\tanh w-g_{w}\left(x\right)}\right)^{2}}\left(\tanh w-g_{w}\left(x\right)\right)^{2}}=$
$=\frac{\tanh w\sqrt{1-\tanh^{2}w}g_{w}^{\prime}\left(x\right)}{\left(\tanh w-g_{w}\left(x\right)\right)\sqrt{\left(\tanh w-g_{w}\left(x\right)\right)^{2}-\left(1-\tanh^{2}w\right)g_{w}^{2}\left(x\right)}}=$
$=\frac{-\tanh w\sqrt{1-\tanh^{2}w}\sin x}{r_{w}\left(x\right)\sqrt{\left(\tanh w-g_{w}\left(x\right)\right)^{2}-\left(1-\tanh^{2}w\right)g_{w}^{2}\left(x\right)}}=$
$=\frac{-\sqrt{2}\tanh w\sqrt{1-\tanh^{2}w}\sin x}{r_{w}\left(x\right)\sqrt{\left(1-\cos x\right)\left(-\left(1+\cos x\right)+2\tanh^{2}w+r_{w}\left(x\right)\right)}}.$
Now we calculate the second derivative. We substitute $r_{w}^{\prime}\left(x\right)=\frac{-\sin x}{r_{w}\left(x\right)}\left(1+\cos x-2\tanh^{2}w\right)$ to obtain the following.
$F_{w}^{\prime\prime}\left(x\right)=\left(-\sqrt{2}\tanh w\sqrt{1-\tanh^{2}w}\right)\cdot$
$\frac{\cos x\cdot r_{w}\left(x\right)\sqrt{1-\cos x}+\sin^{2}x\frac{1+\cos x-2\tanh^{2}w}{r_{w}\left(x\right)}\sqrt{1-\cos x}-\sin^{2}x\frac{r_{w}\left(x\right)+\left(1-\cos x\right)}{2\sqrt{1-\cos x}}}{r_{w}^{2}\left(x\right)\left(1-\cos x\right)\left(-1-\cos x+2\tanh^{2}w+r_{w}\left(x\right)\right)}=$
$=\left(-\frac{\sqrt{2}}{2}\tanh w\sqrt{1-\tanh^{2}w}\right)\cdot\sqrt{\frac{1-\cos x}{-1-\cos x+2\tanh^{2}w+r_{w}\left(x\right)}}\cdot$
$\left(\left(1+\cos x\right)^{2}-4\tanh^{2}w-\left(1+\cos x\right)r_{w}\left(x\right)\right).$
We observe that
$\left(1+\cos x\right)^{2}-4\tanh^{2}w-\left(1+\cos x\right)r_{w}\left(x\right)=r_{w}\left(x\right)\left(r_{w}\left(x\right)-1-\cos x\right)+4\tanh^{2}w\left(\cos x-1\right)<0,$
and hence $-F_{w}$ is strictly concave in the interval $\left(0,\max_{i}\varphi_{i}\right]$. Finally, Jensen’s inequality concludes the proof.
Corollary 4.3.
The area of a regular $\left(2k+1\right)$-gon of a fixed thickness is a strictly increasing monotonic function of $k$. In particular, all ordinary reduced polygons have a smaller area than the circle of the same thickness.
Proof.
If $R_{n}$ is an ordinary reduced polygon of thickness $w$ with $n$ vertices, and $\widetilde{R}_{n}$ denotes a regular $n$-gon of thickness $w$, then by Theorem 4.2,
$\operatorname{\mathrm{area}}\left(R_{n}\right)\leq\operatorname{\mathrm{area}}\left(\widetilde{R}_{n}\right)=\left(n-2\right)\pi-2nF_{w}\left(\frac{\pi}{n}\right)=2n\left(\frac{\pi}{2}-f_{w}\left(g_{w}\left(\frac{\pi}{n}\right)\right)\right)-2\pi=$
$=2n\arccos\frac{g_{w}\left(x\right)\sqrt{1-\tanh^{2}w}}{\tanh w-g_{w}\left(x\right)}-2\pi.$
From (1) and (2) we get
$n=\frac{\pi}{\arccos\frac{g_{w}\left(\frac{\pi}{n}\right)\left(1-\tanh w\cdot g_{w}\left(\frac{\pi}{n}\right)\right)}{\tanh w-g_{w}\left(\frac{\pi}{n}\right)}},$
and hence
$\operatorname{\mathrm{area}}\left(\widetilde{R}_{n}\right)=2\pi\frac{\arccos\frac{g_{w}\left(\frac{\pi}{n}\right)\sqrt{1-\tanh^{2}w}}{\tanh w-g_{w}\left(\frac{\pi}{n}\right)}}{\arccos\frac{g_{w}\left(\frac{\pi}{n}\right)\left(1-\tanh wg_{w}\left(\frac{\pi}{n}\right)\right)}{\tanh w-g_{w}\left(\frac{\pi}{n}\right)}}-2\pi.$
We observe that
$g_{w}^{\prime}\left(x\right)=\frac{-\sin x}{r_{w}\left(x\right)}\left(\tanh w-g_{w}\left(x\right)\right)<0$
for $\tanh w>g_{w}\left(x\right)$, so $g_{w}\left(\frac{\pi}{n}\right)$ increases in $n$. Hence, to see that the area of $\widetilde{R}_{n}$ increases in $n$, it is sufficient to see that the function
$h_{w}\left(x\right)=\frac{\arccos\frac{x\sqrt{1-\tanh^{2}w}}{\tanh w-x}}{\arccos\frac{x\left(1-x\tanh w\right)}{\tanh w-x}}$
increases for $0<x<g_{w}\left(0\right)=\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}$. We note that
$h_{w}^{\prime}\left(x\right)=\frac{\tanh w\cdot\overline{h}_{w}\left(x\right)}{\left(\arccos\frac{x\left(1-x\tanh w\right)}{\tanh w-x}\right)^{2}\left(\tanh w-x\right)\sqrt{\left(1-x^{2}\right)\left(\tanh^{2}wx^{2}+\tanh^{2}w-2x\tanh w\right)}}$
@@
$\overline{h}_{w}\left(x\right)=\left(x^{2}+1-2x\tanh w\right)\arccos\frac{x\sqrt{1-\tanh^{2}w}}{\tanh w-x}-$
$\sqrt{\left(1-\tanh^{2}w\right)\left(1-x^{2}\right)}\arccos\frac{x\left(1-x\tanh w\right)}{\tanh w-x}.$
We note that since $0<x<\tanh w$, the sign of $h_{w}^{\prime}$ and $\overline{h}_{w}$ coincide. Differentiating $\overline{h}_{w}$, we get
$\overline{h}^{\prime}_{w}\left(x\right)=\frac{x\sqrt{1-\tanh^{2}w}}{\sqrt{1-x^{2}}}\arccos\frac{x\left(1-x\tanh w\right)}{\tanh w-x}-2\left(\tanh w-x\right)\arccos\frac{x\sqrt{1-\tanh^{2}w}}{\tanh w-x}.$
Observe that
$\overline{h}_{w}\left(\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}\right)=\overline{h}^{\prime}_{w}\left(\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}\right)=0,$
so to verify that $\operatorname{\mathrm{area}}\left(\widetilde{R}_{n}\right)$ increases in $n$, it is enough to see that $\overline{h}^{\prime\prime}_{w}\left(x\right)>0$, because then for $0<x<\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}$, we deduce $\overline{h}^{\prime}_{w}\left(x\right)<\overline{h}^{\prime}_{w}\left(\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}\right)=0$, and hence $\overline{h}_{w}\left(x\right)>\overline{h}_{w}\left(\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}\right)=0$. Indeed, we have
$\overline{h}^{\prime\prime}_{w}\left(x\right)=2\arccos\frac{x\sqrt{1-\tanh^{2}w}}{\tanh w-x}+\frac{\sqrt{1-\tanh^{2}w}}{\left(1-x^{2}\right)^{\frac{3}{2}}}\arccos\frac{x\left(1-x\tanh w\right)}{\tanh w-x}+$
$\frac{\tanh w\sqrt{1-\tanh^{2}w}}{\left(1-x^{2}\right)\left(\tanh w\right)\sqrt{\tanh^{2}wx^{2}+\tanh^{2}w-2x}}\left(x^{3}-3x+2\tanh w\right),$
and that is positive, as $x^{3}-3x+2\tanh w>0$ if $0<x<\frac{1-\sqrt{1-\tanh^{2}w}}{\tanh w}$.
Finally, we observe that
$\operatorname{\mathrm{area}}\left(R_{n}\right)<\lim_{m\to\infty}\operatorname{\mathrm{area}}\left(\widetilde{R}_{m}\right)=2\pi\left(\cosh\frac{w}{2}-1\right),$
that is the area of a circular disk of radius $\frac{w}{2}$. ∎
As a final remark, we note that the area of the quarter of the disk (which is also reduced) is greater than the area of the circle of the same thickness, and the author has no knowledge of any reduced body that has greater area in $H^{2}$.